# Addendum to “Power-law decay of the fraction of the mixed eigenstates in kicked top model with mixed-type classical phase space”

### v2

@@ IV.1 The overlap index and Wehrl ELM
From the definition of Wehrl ELM following Eq. (18), regular eigenstates characterized by an overlap index of $M=-1$ and condensing on invariant tori are localized, yielding lower values of $\mathcal{L}$. In contrast, chaotic eigenstates with $M=+1$ can be either strongly localized or extended. Previous studies have numerically shown that the probability distribution $P(\mathcal{L})$ of chaotic eigenstates indeed follows a beta distribution [60, 61, 62]. As the system approaches the semiclassical limit, it tends towards a delta distribution $\delta(\mathcal{L}-\langle\mathcal{L}\rangle_{\mathcal{R}})$. It is important to note that in mixed-type systems, to verify this distribution, the Wehrl ELM of the chaotic eigenstates should be further renormalized by the relative area $\mu_{c}$ of the chaotic component in the phase space. In addition to strictly regular and chaotic states, there would exist mixed eigenstates with $|M|<1$ in mixed-type systems, exhibiting varying degrees of localization. This is illustrated in Fig. 7(a), plotting the joint distribution $(\mathcal{L},M)$ for KT at $\gamma=2.6$, alongside selected representative Husimi functions of eigenstates from different regions of the parameter space, as shown in Fig. 7(a1)-(d3). Moreover, Fig. 7(b) presents the SALI plot from the corresponding classical dynamics at $\gamma=2.6$ in phase space, for comparison. At the border of the regular islands and chaotic sea, visible stickiness are denoted by intermediate colors.
@@
Fig. 7 shows that states predominantly cluster at both ends of the overlap index, one end representing regular states and the other chaotic states, while mixed states sparsely distribute in between. This clustering is more evident for QKT at $\gamma=3$, as shown in Fig. 13 (see Appendix C the gallery of states), where larger values of $\gamma$ correspond to increased $\mu_{c}$ – the fraction of chaotic components, and a simpler structure of classical phase space. Examining Husimi functions of the chaotic eigenstates plotted in Fig. 7(a1)-(a3), alongside the corresponding SALI plot, a reduction in the area of the darkest blue regions (with $Q_{n}\leq 10^{-16}$) belonging to regular islands can be observed. The most extended state shown in Fig. 7(a3), infiltrates into the regular islands due to quantum tunneling, and exhibits a minor overlap with the outer tori of the regular islands. The state with the maximum overlap index, as plotted in Fig. 7(a1), is strongly localized around and avoids the outer tori of the regular region. Similar observation extends to another layer of overlap index $M\simeq 0.5$ for states illustrated in Fig. 7(b1)-(b3), and to $M\simeq 0$ in Fig. 7(c1)-(c3), except that the most localized ones now live in the vicinity of the regular islands, with smaller contribution from the chaotic sea. Fig. 7(d1)-(d3) illustrate states from the bottom layer of overlap index. Particularly, Fig. 7(d1) shows a state predominantly localized at a chain of regular islands, which exhibits tunneling between these regular islands. In Fig. 7(d2), the state also resides in the vicinity of the regular island, with visible stickiness surrounding it, appearing more extended. The most localized state shown in Fig. 7(d3) is located exactly at the fixed points.
@@ IV.2 The fraction of mixed eigenstates
In the semiclassical limit, it is expected that mixed states gradually disappear in accordance with the Berry-Robnik picture and PUSC. This is supported by the histograms depicted in Fig. 8, which illustrate the distribution $P(M)$ of the overlap index $M$ as the system size $j$ increases, at different kicking strengths. As the semiclassical limit is approached with increasing $j$, we observe that: (i) Increasing values of $j$ lead to a more pronounced alignment of eigenstates with either regular or chaotic clusters. Specifically, at $\gamma=2.3$, the kicking strength closer to the integrability breaking at $\gamma\simeq 2$, the chaotic cluster centers around $M\simeq 0.8$, correlating with a much more complex phase space structure (see Fig. 12 in Appendix C) and the possible classical partial transport barriers [63, 64, 65, 66, 67, 68]. (ii) Fluctuations among intermediate values of $M$ decrease as the system size $j$ increases. (iii) Increasing $\gamma$ enhances chaos and $\mu_{c}$, reflected in $P(M)$ with a higher peak at $M=+1$ and a lower peak at $M=-1$, and reduced fluctuations for $|M|<1$.
@@
The decay exponent $\zeta$ certainly depends on the interval we consider for $M$. The interval should not be too small, lacking sufficient states for good statistics, or too large to show little variation across different intervals. In Fig. 10, we illustrate the variation of the decay exponent when considering small intervals $M\in[M-\Delta M/2,M+\Delta M/2]$, with $\Delta M=0.3$, for different $\gamma$. The inset displays the goodness of the best-fitting, indicated by $R^{2}$, the coefficient of determination. Disregarding a few imperfect fittings, the data suggests a decay exponent range of $0.450.2<\zeta<$ 0.45. Given the similar power-law decay observed in the fraction of mixed eigenstates in both billiards [23] and quantum Henon-Heilés system [62], along with the Dicke model [69], we conjecture that this behavior is a universal property of mixed states, as the system approaches the semiclassical limit. A more profound open question is whether there exists an upper bound of the decay exponent.
@@
We emphasize that in Ref. [22], the hierarchical states were introduced as a new class of eigenstates in mixed-type systems, distinct from the regular and chaotic ones. These states, observed in the quantum kicked rotor with much simpler phase space structure, predominantly live near the regular islands, with only a minor contribution in the main part of the chaotic sea. They also demonstrate a power-law decay $\hbar^{\alpha}\ (\alpha>0)$, attributed to the transport properties associated with stickiness between regular islands and the chaotic sea. From this definition, the hierarchical states mostly are highly localized states as shown in Fig. 7(b1), (c1), (d1) (or the same labeled states in Fig. 12-13). While the overlap index captures all the mixed states across different degrees of localization, it is influenced also by the flooding of chaotic sea to regular islands or the tunneling between regular islands, not directly related to the effects of classical stickiness. So it would be interesting to study the behavior of mixed states in systems without classical stickiness, like certain lemon billiards [70, 71].
@@ V Conclusions
In this work, we employed the Krylov subspace technique to generate spin coherent states, thereby extending the investigation of the statistics of Husimi functions of eigenstates in the kicked top, a prototype model for studying quantum chaos. Notably, we explored system sizes up to $j=2^{13}$, a significant advancement compared to previous studies of Husimi functions limited to $j$ in the range of a few hundred. We examined the transition to chaos in both classical and quantum case. All the indicators show a good quantum-classical correspondence, specifically the mean Wehrl entropy localization measure in the statistics of eigenstates. With this new spin coherent states generating method, we also demonstrated that the mean Wehrl entropy localization measure of the fully chaotic quantum kicked top, which follows the COE spectral statistics, approaches the CUE value as the system size $j$ increases, which is surprising [50].
More importantly, in the case of a mixed-type kicked top where regular islands and chaotic sea coexist in the classical compact phase space, we examined mixed eigenstates using a joint distribution of the overlap index $M$ and Wehrl localization measure $\mathcal{L}$. $M=+1$ indicates chaotic states, while $M=-1$ indicates regular ones. We extended the study of mixed eigenstates with $|M|<1$ in our previous investigation [26] from $j=400$ to $j=2^{13}$, confirming the power-law decay of the fraction of mixed states across two orders of magnitude in $j$. This provides supporting evidence for the PUSC and the Berry-Robnik picture in the semiclassical limit. Based on this result from quantum kicked top, and similar results from other models [23, 62, 69], we conjecture that the power-law decay is a universal property for mixed eigenstates. This hypothesis merits further theoretical scrutiny.
@@ Appendix A Parity and eigenvector statistics
Numerical result in Fig. 3 shows that $\langle L\rangle$ of $J_{x,-}$ in $|j,m,-\rangle$ basis is close to the COE value. As we have proven, it is the same as the $\langle L\rangle$ of $J_{x}$ on $|j,m\rangle$ basis if $j\gg 1$
@@
The parity of the eigenvector is more clearly reflected in phase space through the spin coherent state $|\theta,\phi\rangle$, that
@@ Appendix B Mean ELM in integrable QKT
In Fig. 3 we have shown that the mean Wehrl ELM $\langle\mathcal{L}\rangle$ of KT maintains a value equal to the mean Wehrl ELM of the unperturbed Hamiltonian $\langle\mathcal{L}\rangle_{J_{x}}$, for $\gamma\leq 2$, which is almost integrable. Fig. 11 shows that at $\gamma=1$, the mean Wehrl entropy $\langle S\rangle\sim\frac{1}{2}\ln N$ and accordingly in the inset $\langle\mathcal{L}\rangle\sim N^{-1/2}$. When contrasting with the uniformly chaotic systems where $\langle S\rangle\sim\ln N$ and $\langle\mathcal{L}\rangle\sim 1$, the distinction between integrable and chaotic cases becomes apparent (see Fig. 5 for the behavior of $\langle S\rangle$ and $\langle\mathcal{L}\rangle$ at $\gamma=8$). In integrable regimes, eigenstates tend to cover a thin band around a circle on the sphere, whereas in chaotic regimes, eigenstates exhibit delocalization across the entire sphere [24].
Figure 11: Mean Wehrl entropy $\langle S\rangle$, versus $N=2j+1$ for QKT with $\gamma=1$ and $\alpha=11\pi/19$. The inset is for mean Wehrl ELM $\langle\mathcal{L}\rangle$. The dashed lines indicate the best fittings.
@@
$|j,$ $m\rangle_{x}=\exp(-i\pi J_{y}/2)|j,m\rangle$
$=R(-\pi/2,0)|j,m\rangle=\sum_{m^{\prime}}d_{m^{\prime}m}^{j}(\pi/2)|j,m^{\prime}\rangle,$ (37)
where $R(\theta,\phi)$ is given in Eq. (11a), $d_{mm^{\prime}}^{j}(\theta)$ is the Wigner $d$-matrix. Moreover, $R(\pi,0)|j,m\rangle=|j,-m\rangle$ while $R(\pi,0)K$ is the conventional time-reversal operator and $K$ is the complex conjugation on Dicke basis. The summation of the mean $q$-moment of Husimi functions of all the eigenvectors of $J_{x}$ can then be written as
@@ Appendix C Gallery of states
Figure 13: Analogous data as in Fig. 7 and 12, but now for $\gamma=3$. Compared with $\gamma=2.3$, in this case, the fixed point at $(\theta,\phi)=(\pi/2,0)$ becomes unstable, which is clearly shown in the SALI plot. Interestingly, the most localized state (d3) occupies identical fixed points, akin to when $\gamma=2.6$.

### v3

@@ IV.1 The overlap index and Wehrl ELM
From the definition of Wehrl ELM following Eq. (20), regular eigenstates characterized by an overlap index of $M=-1$ and condensing on invariant tori are localized, yielding lower values of $\mathcal{L}$. In contrast, chaotic eigenstates with $M=+1$ can be either strongly localized or extended. Previous studies have numerically shown that the probability distribution $P(\mathcal{L})$ of chaotic eigenstates indeed follows a beta distribution [66, 67, 68]. As the system approaches the semiclassical limit, it tends towards a delta distribution $\delta(\mathcal{L}-\langle\mathcal{L}\rangle_{\mathcal{R}})$. It is important to note that in mixed-type systems, to verify this distribution, the Wehrl ELM of the chaotic eigenstates should be further renormalized by the relative area $\mu_{c}$ of the chaotic component in the phase space. In addition to strictly regular and chaotic states, there would exist mixed eigenstates with $|M|<1$ in mixed-type systems, exhibiting varying degrees of localization. This is illustrated in Fig. 7(a), plotting the joint distribution $(\mathcal{L},M)$ for QKT at $\gamma=2.6$, alongside selected representative Husimi functions of eigenstates from different regions of the parameter space, as shown in Fig. 7(a1)-(d3). Moreover, Fig. 7(b) presents the SALI plot from the corresponding classical dynamics at $\gamma=2.6$ in phase space, for comparison. At the border of the regular islands and chaotic sea, visible stickiness are denoted by intermediate colors.
@@
Fig. 7(a) shows that states predominantly cluster at both ends of the overlap index, one end representing regular states and the other chaotic states, while mixed states sparsely distribute in between. In comparison, for smaller kicking strengths, such as $\gamma=2.3$, shown in Fig. 13(a) (see Appendix C the gallery of states), the clustering is less pronounced. At $\gamma=3$ (Fig. 14(a)), the clustering becomes more evident, as larger values of $\gamma$ correspond to larger $\mu_{c}$ – the fraction of chaotic components, and a simpler structure of classical phase space, resulting in a smaller region between the chaotic sea and regular islands where stickiness occurs. Examining Husimi functions of the chaotic eigenstates plotted in Fig. 7(a1)-(a3), alongside the corresponding SALI plot, a reduction in the area of the darkest blue regions (with $Q_{n}\leq 10^{-16}$) belonging to regular islands can be observed. The most extended state shown in Fig. 7(a3), infiltrates into the regular islands due to quantum tunneling, and exhibits a minor overlap with the outer tori of the regular islands. The state with the maximum overlap index, as plotted in Fig. 7(a1), is strongly localized around and avoids the outer tori of the regular region. Similar observation extends to another layer of overlap index $M\simeq 0.5$ for states illustrated in Fig. 7(b1)-(b3), and to $M\simeq 0$ in Fig. 7(c1)-(c3), except that the most localized ones now live in the vicinity of the regular islands, with smaller contribution from the chaotic sea. Fig. 7(d1)-(d3) illustrate states from the bottom layer of overlap index. Particularly, Fig. 7(d1) shows a state predominantly localized at a chain of regular islands, which exhibits tunneling between these regular islands. In Fig. 7(d2), the state also resides in the vicinity of the regular island, with visible stickiness surrounding it, appearing more extended. The most localized state shown in Fig. 7(d3) is located exactly at the fixed points.
@@ IV.2 The fraction of mixed eigenstates
In the semiclassical limit, it is expected that mixed states gradually disappear in accordance with the Berry-Robnik picture and PUSC. This is supported by the histograms depicted in Fig. 8, which illustrate the distribution $P(M)$ of the overlap index $M$ as the system size $j$ increases, at different kicking strengths. As the semiclassical limit is approached with increasing $j$, we observe that: (i) Increasing values of $j$ lead to a more pronounced alignment of eigenstates with either regular or chaotic clusters. Specifically, at $\gamma=2.3$, the kicking strength closer to the integrability breaking at $\gamma\simeq 2$, the chaotic cluster centers around $M\simeq 0.8$, correlating with a much more complex phase space structure (see Fig. 13 in Appendix C) and the possible classical partial transport barriers [69, 70, 71, 72, 73, 37]. (ii) Fluctuations among intermediate values of $M$ decrease as the system size $j$ increases. (iii) Increasing $\gamma$ enhances chaos and $\mu_{c}$, reflected in $P(M)$ with a higher peak at $M=+1$ and a lower peak at $M=-1$.
@@
The decay exponent $\zeta$ certainly depends on the interval we consider for $M$. The interval should not be too small, lacking sufficient states for good statistics, or too large to show little variation across different intervals. In Fig. 10, we illustrate the variation of the decay exponent when considering small intervals $M\in[M-\Delta M/2,M+\Delta M/2]$, with $\Delta M=0.3$, for different $\gamma$. The inset displays the goodness of the best-fitting, indicated by $R^{2}$, the coefficient of determination. Disregarding a few imperfect fittings, the data suggests a decay exponent range of $0.450.2<\zeta<$ 0.45. Given the similar power-law decay observed in the fraction of mixed eigenstates in both billiards [23] and quantum Hénon-Heiles system [68], along with the Dicke model [74], we conjecture that this behavior is a universal property of mixed states, as the system approaches the semiclassical limit. A more profound open question is whether there exists an upper bound of the decay exponent.
We emphasize that in Ref. [22], the hierarchical states were introduced as a new class of eigenstates in mixed-type systems, distinct from the regular and chaotic ones. These states, observed in the quantum kicked rotor with much simpler phase space structure, predominantly live near the regular islands, with only a minor contribution in the main part of the chaotic sea. They also demonstrate a power-law decay $\hbar^{\alpha}\ (\alpha>0)$, attributed to the transport properties associated with stickiness between regular islands and the chaotic sea. From this definition, the hierarchical states mostly are highly localized states as shown in Fig. 7(b1), (c1), (d1) (or the same labeled states in Fig. 13-14). While the overlap index captures all the mixed states across different degrees of localization, it is influenced also by the flooding of chaotic sea to regular islands or the tunneling between regular islands, not directly related to the effects of classical stickiness. So it would be interesting to study the behavior of mixed states in systems without classical stickiness, like certain lemon billiards [75, 76].
@@ V Conclusions
In this work, we employed the Krylov subspace technique to generate spin coherent states, thereby extending the investigation of the statistics of Husimi functions of eigenstates in the kicked top, a prototype model for studying quantum chaos. Notably, we explored system sizes up to $j=2^{13}$, a significant advancement compared to previous studies of Husimi functions limited to $j$ in the range of a few hundred. We examined the transition to chaos in both classical and quantum case. All the indicators show a good quantum-classical correspondence, specifically the mean Wehrl entropy localization measure in the statistics of eigenstates. With this new spin coherent states generating method, we also demonstrated that the mean Wehrl entropy localization measure of the fully chaotic quantum kicked top, which follows the COE spectral statistics, approaches the CUE value as the system size $j$ increases, which is surprising [53]. This indicates that the statistics followed by the eigenvalues and eigenvectors do not necessarily follow the same ensemble. We proved that coherent states located on symmetry lines of the classical phase space exhibit COE eigenvector statistics but have a zero measure, while other coherent states display eigenvector statistics consistent with CUE.
More importantly, in the case of a mixed-type kicked top where regular islands and chaotic sea coexist in the classical compact phase space, we examined mixed eigenstates using a joint distribution of the overlap index $M$ and Wehrl localization measure $\mathcal{L}$. $M=+1$ indicates chaotic states, while $M=-1$ indicates regular ones. We extended the study of mixed eigenstates with $|M|<1$ in our previous investigation [26] from $j=400$ to $j=2^{13}$, confirming the power-law decay of the fraction of mixed states across two orders of magnitude in $j$. This provides supporting evidence for the PUSC and the Berry-Robnik picture in the semiclassical limit. Based on this result from quantum kicked top, and similar results from other models [23, 68, 74], we conjecture that the power-law decay is a universal property for mixed eigenstates. This hypothesis, particularly the connection with classical stickiness or partial barriers, merits further theoretical scrutiny.
@@ Appendix A Parity and eigenvector statistics
Numerical result in Fig. 4 shows that $\langle L\rangle$ of $J_{x,-}$ in the basis $|j,m,-\rangle$ is close to the COE value. As we have proven, it is the same as the $\langle L\rangle$ of $J_{x}$ on $|j,m\rangle$ basis if $j\gg 1$
@@
The parity of the eigenvector is more clearly reflected in phase space through the spin coherent state $|\theta,\phi\rangle$,
@@ Appendix B Mean ELM in integrable QKT
In Fig. 4 we have shown that the mean Wehrl ELM $\langle\mathcal{L}\rangle$ of KT maintains a value equal to the mean Wehrl ELM of the unperturbed Hamiltonian $\langle\mathcal{L}\rangle_{J_{x}}$, for $\gamma\leq 2$, which is almost integrable. Fig. 11 shows that at $\gamma=1$, the mean Wehrl entropy $\langle S\rangle\sim\frac{1}{2}\ln N$ and accordingly in the inset $\langle\mathcal{L}\rangle\sim N^{-1/2}$. When contrasting with the uniformly chaotic systems where $\langle S\rangle\sim\ln N$ and $\langle\mathcal{L}\rangle\sim 1$, the distinction between integrable and chaotic cases becomes apparent (see Fig. 5 for the behavior of $\langle S\rangle$ and $\langle\mathcal{L}\rangle$ at $\gamma=8$). In integrable regimes, eigenstates tend to cover a thin band around a circle on the sphere, whereas in chaotic regimes, eigenstates exhibit delocalization across the entire sphere [24]. We have examined $\langle\mathcal{L}\rangle$ with increasing number of grid points to $N_{p}=250\times 500$, for $j=2^{12}$. In Fig. 12 it shows that the mean Wehrl ELM is nearly unchanged versus the increasing number of grid points, for three different $\gamma$.
Figure 11: Mean Wehrl entropy $\langle S\rangle$, versus $N=2j+1$ for QKT with $\gamma=1$ and $\alpha=11\pi/19$. The inset is for mean Wehrl ELM $\langle\mathcal{L}\rangle$, calculated from $8\times 10^{4}$ coherent states held on $200\times 400$ grid points in phase space. The dashed lines indicate the best fittings.
@@
$|j,m\rangle_{x}=R(-\frac{\pi}{2},0)|j,m\rangle=\sum_{m^{\prime}}d_{m^{\prime}m}^{j}(\pi/2)|j,m^{\prime}\rangle,$ (40)
where $R(\theta,\phi)$ is given in Eq. (13a), $d_{mm^{\prime}}^{j}(\theta)$ is the Wigner $d$-matrix. The summation of the mean $q$-moment of Husimi functions of all the eigenvectors of $J_{x}$ can then be written as
@@ Appendix C Gallery of states
Figure 14: Analogous data as in Fig. 7 and 13, but now for $\gamma=3$. Compared with $\gamma=2.3$, in this case, the fixed point at $(\theta,\phi)=(\pi/2,0)$ becomes unstable, which is clearly shown in the SALI plot. Interestingly, the most localized state (d3) occupies identical fixed points, akin to when $\gamma=2.6$.